# Modeling dissipation in quantum active matter

### v1

@@ II.3 Numerical procedure
where $\overline{\cdots}$ denotes the average over prescribed trajectories $x_{c}(t)$. In (14), we first evaluate the trace of the observable (i.e., perform the quantum averaging for given instantaneous $x_{c}(t)$), and only afterwards average over the prescribed trajectories $x_{c}(t)$. This ordering is natural since $x_{c}(t)$ is treated as an externally fixed input, not as part of the quantum dynamics itself. However, the averages commute: Eq. (4) involves purely classical sampling, so no interference effects arise, and the order of averaging is irrelevant.
@@
II.4 Potential experimental realization
A possible realization of the model considered here is a quantum-mechanical particle in a trap centered at position $x_{c}(t)$. With a computer, the trajectory of the center $\{x_{c}(t)\}$ is generated, which is possible since the dynamics of $x_{c}$ does not depend on $\hat{\rho}$. By moving the center of the potential, we can realize the Hamiltonian (2). At the final time, the spatial probability distribution of the quantum particle is obtained from the wavefunction. By repeating this procedure for a set of $\{x_{c}(t)\}$, the average value of the quantum mean squared displacement, explained in the following subsection, can be measured. A sketch of such experimental realization is shown in Fig. 1.
Figure 1: Schematic representation of a possible experiment – an external parabolic potential (yellow) captures a quantum particle (blue wave function). The potential moves along a trajectory $x_{c}(t)$ corresponding to colored noise. Since the quantum particle is constantly driven away from the instantaneous trap center, the system never truly reaches equilibrium, and the quantum particle thereby mimics the properties of active motion.
III Results
In Fig. 2, we compare MSDs for different dissipation strengths of the static Lindblad master equation (II.2.1), using the MSD of the driving protocol, $\overline{x_{c}^{2}(t)}$, as a reference. Specifically, $x_{c}(t)$ in (1) corresponds to a passive inertial particle, exhibiting $t^{3}$-scaling at short times before crossing over to diffusion [44]. When $x_{c}(t)$ acts as a colored-noise input to the moving potential, it drives the particle in a way that breaks the detailed balance, allowing us to study how quantum dissipators influence motion that mimics active dynamics.
For weak dissipation, the MSD exhibits a scaling close to $t^{6}$, as theoretically predicted for the non-dissipative case [5]. However, when the dissipation increases, the effect of the dissipator $\hat{\mathcal{D}}$ becomes more significant, driving the particle towards the origin and thereby reducing the effectiveness of the trapping potential in controlling the quantum particle.
A convenient way to understand this effect is to describe the system via the Wigner quasi-probability function $W(x,p)$. The Wigner function [45] allows one to represent the state of a quantum system by distribution in position-momentum space that is related to the probability distribution in classical statistical mechanics [46], but can also be negative. This allows, among other things, to write quantum-mechanical equations in a way that is very analogous to dynamical equations commonly studied in soft matter physics [47, 48], which makes them a promising tool in the study of quantum active matter. In the Wigner representation, the quantum master equation (II.1) corresponds to a Fokker-Planck equation for $W(x,p)$, providing an alternative but fully equivalent description of the same quantum dynamics (see Appendix B for the derivation). This Fokker-Planck equation reads
Figure 2: MSD of quantum-active particle with static Lindblad dissipator Eq. (II.2.1) for weak, intermediate and strong dissipation respectively. The parameters of $x_{c}$ are $D=0.01,\tau=10$, and the corresponding classical $\operatorname*{MSD}=\overline{x_{c}^{2}(t)}$ of these trajectories is shown as a black line. The dissipation strength $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is set by $\nu_{+}=10^{-8}$ and $\nu_{-}=10^{-2}$ (weak), $\nu_{-}=10^{0}$ (intermediate), $\nu_{-}=10^{1}$ (strong).
$\partial_{t}W\!$ $=$ $\!\partial_{i}(f_{i}W)+\hbar\partial^{2}_{ij}(g_{ij}W),$ (15a)
$f_{i}\!$ $=$ $\!\begin{pmatrix}-\frac{p}{m}+\frac{\gamma}{4}x\\ m\omega^{2}(x-x_{c})+\frac{\gamma}{4}p\end{pmatrix}$ (15b)
$g_{ij}\!$ $=$ $\!\begin{pmatrix}\frac{\gamma}{8m\omega}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)&0\\ 0&\frac{\gamma m\omega}{8}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\end{pmatrix}$.
Here and in what follows the indices $i$ and $j$ represent coordinates $x$ and $p$. We note that the Wigner Fokker-Planck equation contains terms that are not present in a classical dissipative harmonic oscillator,
$m\ddot{x}(t)+\gamma\dot{x}(t)=-m\omega^{2}(x(t)-x_{c})+\sqrt{\gamma k_{B}T}\eta(t),$ (16)
@@ III Results
for which the Fokker-Planck equation reads [49]:
$\partial_{t}P\!$ $=$ $\!\partial_{i}(f_{i}P)+\partial^{2}_{ij}(g_{ij}P),$ (17a)
$f_{i}\!$ $=$ $\!\begin{pmatrix}-\frac{p}{m}\\ m\omega^{2}(x-x_{c})+\gamma p\end{pmatrix}$, (17b)
$g_{ij}\!$ $=$ $\!\begin{pmatrix}0&0\\ 0&\gamma k_{B}T\end{pmatrix}$.
To make the extremum points of the steady-state probability distribution explicit, here we fix $x_{c}$ so that the correct quantum particle’s probability distribution is centered at $x_{c}$ in the steady state. In the full system, however, $x_{c}=x_{c}(t)$ evolves as a colored noise; consequently, the particle is constantly driven away from the instantaneous trap center, and, as is mentioned before, the system never truly reaches equilibrium. Hence, the dynamics is genuinely active.
To illustrate the differences for different dissipation strengths, in Fig. 3 we present the instantaneous steady states of their corresponding Fokker-Planck equations $W(x,p)$ for fixed $x_{c}$. For Eq. (15), it is given by a Gaussian state
@@
Figure 3: Steady-state Wigner function $W(x,p)$ for static Lindblad dissipator (18). The position of the trapping potential is fixed at $x_{c}=3$ for (a) weak dissipation and (b) strong dissipation. The amplitude of the Wigner function values is given by the corresponding color representation; the graphs on the left and right correspond to the peak of the Wigner function at the extremum in $x$ and $p$, respectively. The green dashed lines indicate the position ($x=x_{c}$) of the fixed ($p=0$) harmonic potential. The dissipation strength $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is set by $\nu_{+}=10^{-8}$ and $\nu_{-}=10^{-2}$ (weak), $\nu_{-}=10^{1}$ (strong).
This issue is resolved with the translated Lindblad dissipator(II.2.2), where the modified creation and annihilation operators $\hat{\tilde{a}}^{\dagger}(t),\hat{\tilde{a}}(t)$ are adjusted to the position of the potential minimum $x_{c}(t)$. As demonstrated by the MSDs in Fig. 4, this correction makes the quantum-active particle behavior consistent with the long-time MSD behavior of the active protocol $x_{c}(t)$. Moreover, the result is consistent with the classical active particle dynamics, exhibiting a transition in the MSD scaling from $t^{6}$ in the underdamped (weak dissipation) regime to $t^{4}$ in the overdamped (strong dissipation) regime [5].
Figure 4: MSD of quantum-active particle for translated Lindbald dissipator (II.2.2) for weak, intermediate and strong dissipation respectively. The parameters of $x_{c}$ are $D=0.01,\tau=10$, and the corresponding classical $\operatorname*{MSD}=\overline{x_{c}^{2}(t)}$ of these trajectories is shown as a black line. The dissipation strength $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is set by $\nu_{+}=10^{-8}$ and $\nu_{-}=10^{-2}$ (weak), $\nu_{-}=10^{0}$ (intermediate), $\nu_{-}=10^{1}$ (strong).
In terms of the Wigner function, the quantum master equation with dissipator (II.2.2) reads
$\partial_{t}W\!$ $=$ $\!\partial_{i}(f_{i}W)+\hbar\partial^{2}_{ij}(g_{ij}W),$ (20a)
$f_{i}\!$ $=$ $\!\begin{pmatrix}-\frac{p}{m}+\frac{\gamma}{4}(x-x_{c})\\ m\omega(x-x_{c})+\frac{\gamma}{4}p\end{pmatrix}$ (20b)
$g_{ij}\!$ $=$ $\!\begin{pmatrix}\frac{\gamma}{8m\omega}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)&0\\ 0&\frac{\gamma m\omega}{8}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\end{pmatrix}$,
which for fixed $x_{c}$ has the stationary solution
$W(x,p)\!$
$=$ $\mathcal{N}\exp\left[-\tanh\left(\!\frac{\hbar\omega}{k_{B}T}\!\right)\!\left(\frac{m\omega(x\!-\!x_{c})^{2}}{\hbar}\!+\!\frac{p^{2}}{m\hbar\omega}\right)\right],$
ensuring that the quantum particle follows the active protocol $x_{c}$. The steady-state Wigner function, shown in Fig. 5 in this case has the extremum of the distribution at $x=x_{c}$ and $p=0$ for both weak (Fig. 5(a)) and strong (Fig. 5(b)) dissipation strengths.
Figure 5: Steady-state Wigner function $W(x,p)$ for dynamic Lindblad master equation (III). The position of the trapping potential is fixed at $x_{c}=3$ for (a) weak dissipation, $\nu_{+}=10^{-8},\nu_{-}=10^{-2}$; and (b) strong dissipation $\nu_{+}=10^{-8},\nu_{-}=10^{1}$. The amplitude of the Wigner function values is given by the corresponding color representation; the graphs on the left and right correspond to the peak of the Wigner function at the extremum in $x$ and $p$, respectively. The green dashed lines indicate the position ($x=x_{c}$) of the fixed ($p=0$) harmonic potential. The dissipation strength $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is set by $\nu_{+}=10^{-8}$ and $\nu_{-}=10^{-2}$ (weak), $\nu_{-}=10^{1}$ (strong).
Although the translated Lindblad master equation correctly reproduces active motion even under strong dissipation, it does not display classical active motion in the classical limit, since the structure of (20) is different from that of (17) in the limit $\hbar\to 0$, specifically in the force fields $f_{i}$. The classical limit is however recovered for the Agarwal dissipator (II.2.3), for which the Wigner function representation,
$\partial_{t}W$ $=$ $\partial_{i}(f_{i}W)+\hbar\partial^{2}_{ij}(g_{ij}W),$ (22a)
$f_{i}$ $=$ $\begin{pmatrix}-\frac{p}{m}\\ m\omega(x-x_{c})+\frac{\gamma p}{4}\end{pmatrix}$ (22b)
$g_{ij}$ $=$ $\begin{pmatrix}0&0\\ 0&\frac{\gamma m\omega}{8}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\end{pmatrix}$, (22c)
is analogous to the classical dissipative harmonic oscillator (17), with both the force fields $f_{i}$ and diffusion coefficient $g_{ij}$ identical between the classical and quantum cases in the limit $\hbar\to 0$.
@@
Figure 6: MSD of quantum-active particle with Agarwal dissipator (II.2.3) for weak, intermediate and strong dissipation respectively. The parameters of $x_{c}$ are $D=0.01,\tau=10$, and the corresponding classical $\operatorname*{MSD}=\overline{x_{c}^{2}(t)}$ of these trajectories is shown as a black line. The dissipation strength $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is set by $\nu_{+}=10^{-8}$ and $\nu_{-}=10^{-2}$ (weak), $\nu_{-}=10^{0}$ (intermediate), $\nu_{-}=10^{1}$ (strong).
The stationary solution of Eq. (22) for fixed $x_{c}$ is given by Eq. (III) and therefore identical to that of translated Lindblad master equation (II.2.2). However, while both Fokker-Planck equations have the same steady-state for the fixed $x_{c}$, their dynamics differ, leading to further discrepancies under a time-dependent driving protocol $x_{c}(t)$. In the dynamic Lindblad master equation, both the Hamiltonian part and the dissipator $\hat{\mathcal{D}}$ in the evolution equation (II.1) drive the quantum particle toward the minimum of the moving trap $x_{c}(t)$. In contrast, for the Agarwal dissipator, the driving is performed solely by the Hamiltonian part, while the dissipator $\hat{\mathcal{D}}$ acts purely as a source of friction, as is evident from the corresponding Wigner function representations (20), (22).
The differences between Lindblad and Agarwal dynamics are further illustrated by comparing Fig. 4 and Fig. 6, which show the MSDs of the corresponding systems. At intermediate and long time scales, for both dynamic Lindblad and Agarwal dissipators the dynamics of quantum particles follows that of the trapping potential and captures the properties of the active motion, characterized by a quadratic and later linear diffusive MSD regimes [11]. However, the different nature of the considered dissipators leads to distinct MSD regimes at very short and short time scales. In the dynamic Lindblad case, quantum fluctuations induced by the dissipator – generated by the creation and annihilation operators of the Lindblad dissipator – produce a diffusive regime at very short time scales (Fig. 4). By contrast, this short-time diffusive regime is absent for the Agarwal dissipator (Fig. 6): the quantum particle simply follows the stochastic trajectory $x_{c}(t)$ with an inertial delay and has the same $t^{3}$-scaling as the driving protocol. Moreover, for the dynamic Lindblad dissipator (Fig. 4), stronger dissipation enables the system to establish active motion more rapidly than in the weakly dissipative case. For the Agarwal dissipator (Fig. 6), the situation is reversed – the frictional role of the dissipator $\hat{\mathcal{D}}$ delays the onset of active motion when dissipation is strong.
IV Conclusions and outlook
In this work, we addressed the question of how different types of quantum dissipators influence the emergence of active-like dynamics of a quantum particle, aiming to understand how the environment affects the characteristics of motion reminiscent of classical active matter. By comparing static and dynamically modified Lindblad dissipators as well as the Agarwal dissipator, we systematically explored the interplay between dissipation strength, external driving, and quantum effects. For weak dissipation, we find that the quantum particle exhibits active-like behavior for all dissipator models considered. However, in the case of the static Lindblad dissipator, strong dissipation suppresses the particle’s ability to follow the trapping potential, indicating the need for a dynamical correction of the dissipator. With such dynamical modifications, the Lindblad dissipator can reproduce the driving-induced motion across all dissipation regimes. Both the translated Lindblad and Agarwal dissipators yield motion that follows the driving potential. The translated Lindblad formulation satisfies the CPTP requirement, whereas the Agarwal equation yields dynamics consistent with the classical limit.
Our findings suggest several directions for future research. On the experimental side, possible realizations may be found in cold-atom systems, in particular in moving optical traps where the presence of noisy laser intensity is known to induce additional heating mechanisms [50]. From a theoretical perspective, developing strategies for controlling quantum particles in stochastic [51] and dissipative [52] environments will be crucial for bridging the gap between abstract models and experimentally accessible setups.
Acknowledgements
M.t.V. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB 1551, Project-ID 464588647. S.L. thanks R. Kosloff for a discussion on modeling an open quantum system.
Appendix A Translation of the quantum dissipators
@@
We consider the evolution equation (II.1) translated by $x_{c}$, i.e., we apply the translation generator $\hat{T}(x_{c})$,
@@
and the dissipator $\hat{\mathcal{D}}$. In other words, a state vector is redefined ($|\psi\rangle=\hat{T}|\psi\rangle$ and $\hat{\rho}=\hat{T}\hat{\rho}\hat{T}^{\dagger}$), and all operators are transformed as $\hat{T}\hat{b}\hat{T}^{\dagger}$.
For the Lindblad dissipator,
$\hat{\mathcal{D}}_{L}$ $=$ $\nu_{-}[\hat{a}\hat{\rho}(t)\hat{a}^{\dagger}-\frac{1}{2}\left\{\hat{a}^{\dagger}\hat{a},\hat{\rho}(t)\right\}]$ (25)
@@
$=$ $-\frac{i}{\hbar}[\hat{H}(t),\hat{\rho}(t)]+\hat{\mathcal{D}}_{TL}.$ (33)
@@
which clarifies the role of Lindblad terms in Eq. (A) as excitations and de-excitations of the quantum harmonic oscillator centered at $x_{c}(t)$.
For the Agarwal dissipator $\hat{\mathcal{D}}_{A}$ (II.2.3), we have:
@@
all terms containing $x_{c}$ are commutators with a c-number which therefore vanish. Thus, the Agarwal dissipator preserves its form under $\hat{T}(x_{c})$.
Appendix B Wigner-Weyl transform
B.1 Lindblad dissipator
We consider here the creation-annihilation operators in the form
[…]
where $\delta=0$ for $\hat{\mathcal{D}}_{L}$ (7) and $\delta=1$ for $\hat{\mathcal{D}}_{TL}$ (12).
$\left(\partial_{t}\rho\right)_{W}$ $\equiv$ $\partial_{t}W(x,p,t),$ (38a)
where $\star$ is the Moyal product defined as:
or equivalently, expanded in powers of $\hbar$:
${\frac{1}{i\hbar}}(f\star g-g\star f)=\{f,g\}+O(\hbar^{2}).$ (41)
For the Lindblad dissipator,
$\left(\hat{A}\hat{\rho}\hat{B}\right)_{W}$ $=$ $(A_{W}\star W)\star B_{W}$ (42)
Given that operators $\hat{A},\hat{B}$ in the dissipator $\hat{\mathcal{D}}$ are linear in $x,p$, Eq. (43) reduces to:
$\left(\hat{A}\hat{\rho}\hat{B}\right)_{W}$ $=$ $A_{W}WB_{W}$ (44)
Thus, the triple products relevant in our case are
$\left(\hat{a^{\dagger}}\hat{\rho}\hat{a}\right)_{W}$ $=$ $|a|^{2}W$ (45)
$+$ $\frac{i\hbar}{2}\left((\alpha\beta-(-\beta)\alpha)W+a^{\dagger}(\partial_{x}W\beta-\partial_{p}W\alpha)+a(\alpha\partial_{p}W-(-\beta)\partial_{x}W)\right)$
$=$ $|a|^{2}W-\frac{1}{2}\left(W+p\partial_{p}W+(x-\delta x_{c})\partial_{x}W\right)+\frac{m\omega\hbar}{8}\partial_{x}^{2}W+\frac{\hbar}{8m\omega}\partial_{p}^{2}W;$
$\left(\hat{a}\hat{\rho}\hat{a^{\dagger}}\right)_{W}$ $=$ $|a|^{2}W$ (46)
$+$ $\frac{i\hbar}{2}\left((-\alpha\beta-\beta\alpha)W+a(-\partial_{x}W\beta-\partial_{p}W\alpha)+a^{\dagger}(\alpha\partial_{p}W-\beta\partial_{x}W)\right)$
$=$ $|a|^{2}W+\frac{1}{2}\left(W+p\partial_{p}W+(x-\delta x_{c})\partial_{x}W\right)+\frac{m\omega\hbar}{8}\partial_{x}^{2}W+\frac{\hbar}{8m\omega}\partial_{p}^{2}W;$
For the anti-commutators, we obtain
$\frac{1}{2}\left\{\hat{a}\hat{a}^{\dagger},\hat{\rho}\right\}$ $=$ $\left(|a|^{2}+\frac{i\hbar}{2}(\partial_{x}a\partial_{p}a^{\dagger}-\partial_{p}a\partial_{x}a^{\dagger})\right)W-\frac{\hbar^{2}}{8}(\partial_{x}^{2}|a|^{2}\partial_{p}^{2}W-\partial_{p}^{2}|a|^{2}\partial^{2}_{x}W)$ (47)
$=$ $(|a|^{2}+\frac{1}{2})W-\frac{\hbar^{2}}{4}(\alpha^{2}\partial_{p}^{2}W-\beta^{2}\partial^{2}_{x}W),$
$\frac{1}{2}\left\{\hat{a}^{\dagger}\hat{a},\hat{\rho}\right\}$ $=$ $\left(|a|^{2}+\frac{i\hbar}{2}(\partial_{x}a^{\dagger}\partial_{p}a-\partial_{p}a^{\dagger}\partial_{x}a)\right)W-\frac{\hbar^{2}}{8}(\partial_{x}^{2}|a|^{2}\partial_{p}^{2}W-\partial_{p}^{2}|a|^{2}\partial^{2}_{x}W)$ (48)
$=$ $(|a|^{2}-\frac{1}{2})W-\frac{\hbar^{2}}{4}(\alpha^{2}\partial_{p}^{2}W-\beta^{2}\partial^{2}_{x}W).$
Finally, we arrive at
$(\hat{\mathcal{D}})_{W}$ $=$ $\nu_{+}\Bigg(-\frac{1}{2}\left(2W+p\partial_{p}W+(x-\delta x_{c})\partial_{x}W\right)+\frac{m\omega\hbar}{4}\partial_{x}^{2}W+\frac{\hbar}{4m\omega}\partial_{p}^{2}W\Bigg)$ (49)
$\partial_{t}W$ $=$ $\partial_{i}(f_{i}W)+\hbar\partial^{2}_{ij}(g_{ij}W),$ (50a)
$f_{i}$ $=$ $\begin{pmatrix}-\frac{p}{m}+\frac{\gamma}{4}(x-\delta x_{c})\\ m\omega^{2}(x-x_{c})+\frac{\gamma}{4}p\end{pmatrix}$ (50b)
$g_{ij}$ $=$ $\begin{pmatrix}\frac{\gamma}{8m\omega}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)&0\\ 0&\frac{\gamma m\omega}{8}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right).\end{pmatrix}$ (50c)
Here, we have used the following relations:
$\nu_{-}$ $=$ $\frac{\gamma}{2}(\bar{n}+1),\ \nu_{+}=\frac{\gamma}{2}\bar{n},$ (51)
Setting $\delta=0$ (static Lindblad) restores (15), and $\delta=1$ (translated Lindblad) yields (20).
B.2 Agarwal dissipator
[…]
$=$ $\frac{\hbar\gamma}{4}(W+p\partial_{p}W)+\frac{\gamma m\omega}{8}\hbar\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\partial_{p}^{2}W.$
Combining (38) with (61) yields (22).

### v2

@@ A.2 Role of the Lindblad dissipator form
This variant is appropriate for the case of small damping, which was the focus of interest in Ref. [5]. Here, we discuss why for the case of stronger damping one should use the form (11), which can be derived from (A.2) by applying a translation operator (see Appendix B).
@@
The Fokker-Planck equation for the Wigner function corresponding to the dissipator (A.2) reads
$\partial_{t}W\!$ $=$ $\!\sum_{i=x,p}\partial_{i}(f_{i}W)+\sum_{i,j=x,p}\partial^{2}_{ij}(g_{ij}W),$ (37a)
$f_{x}$ $=$ $-\frac{p}{m}+\frac{\gamma}{4}x,\quad f_{p}=m\omega^{2}(x-x_{c})+\frac{\gamma}{4}p$ (37b)
$g_{xx}$ $=$ $\frac{\hbar\gamma}{8m\omega}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right),\quad g_{pp}=\frac{\hbar\gamma m\omega}{8}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right),\quad g_{xp}=g_{px}=0.$
Figure 5: Steady-state Wigner function $W(x,p)$ for “static” Lindblad dissipator (38). The position of the trapping potential is fixed at $x_{c}=3$ for (a) weak dissipation and (b) strong dissipation. The amplitude of the Wigner function values is given by the corresponding color representation; the graphs on the left and right correspond to the peak of the Wigner function at the extremum in $x$ and $p$, respectively. The green dashed lines indicate the position ($x=x_{c}$) of the fixed ($p=0$) harmonic potential. The thermalization rate $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is varied at fixed temperature $T=\hbar\omega/\left(k_{B}\ln\left[\nu_{-}/\nu_{+}\right]\right)$ (9c) with $\nu_{-}/\nu_{+}=100$ for $\nu_{+}=10^{-4}$ (weak) and $\nu_{+}=10^{-1}$ (strong).
For fixed $x_{c}$, the steady-state of Eq. (37) reads:
@@
Figure 6: MSD of quantum-active particle with “static” Lindblad dissipator Eq. (A.2) for weak, intermediate and strong dissipation respectively. The parameters of $x_{c}$ are $D=0.01,\tau=10$, and the corresponding classical $\operatorname*{MSD}=\overline{x_{c}^{2}(t)}$ of these trajectories is shown as a black line. The thermalization rate $\gamma=2(\nu_{-}-\nu_{+})$ (9b) is varied at fixed temperature $T=\hbar\omega/\left(k_{B}\ln\left[\nu_{-}/\nu_{+}\right]\right)$ (9c) with $\nu_{-}/\nu_{+}=100$ for $\nu_{+}=10^{-4}$ (weak), $\nu_{+}=10^{-2}$ (intermediate) and $\nu_{+}=10^{-1}$ (strong).
A.3 Agarwal dissipator
[…]
$=$ $\frac{\hbar\gamma}{4}(W+p\partial_{p}W)+\frac{\hbar\gamma m\omega}{8}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\partial_{p}^{2}W.$
Combining (19) with (47) yields (16).
Appendix B Translation of the quantum dissipators
Here, we show how one can obtain the general Lindblad dissipator (11) by applying the translation generator $\hat{T}(x_{c})$ to (A.2). Physically, this corresponds to taking into account the trap motion when modeling the bath dynamics. This is (also for readers from classical active soft matter physics) worth discussing here explicitly because it marks a contrast to the case of classical baths, which tend to have an Agarwal-like structure and thus do not require to take translations into account for bath models in this way.
@@ Appendix B Translation of the quantum dissipators
We consider the evolution equation (II.1) with the static Lindblad dissipator (A.2) translated by $x_{c}$, i.e., we apply the translation generator $\hat{T}(x_{c})$,
@@
and the dissipator $\mathcal{D}_{SL}$}. In other words, a state vector is redefined ($|\psi\rangle=\hat{T}|\psi\rangle$ and $\hat{\rho}=\hat{T}\hat{\rho}\hat{T}^{\dagger}$), and all operators are transformed as $\hat{T}\hat{b}\hat{T}^{\dagger}$.
For the Lindblad dissipator (A.2),
${\mathcal{D}}_{SL}$ $=$ $\nu_{-}[\hat{a}\hat{\rho}(t)\hat{a}^{\dagger}-\frac{1}{2}\left\{\hat{a}^{\dagger}\hat{a},\hat{\rho}(t)\right\}]$ (50)
@@
$=$ $-\frac{i}{\hbar}[\hat{H}(t),\hat{\rho}(t)]+{\mathcal{D}}_{L}.$ (58)
@@
which clarifies the role of Lindblad terms in Eq. (B) as excitations and de-excitations of the quantum harmonic oscillator centered at $x_{c}(t)$.
For the Agarwal dissipator ${\mathcal{D}}_{A}$ (II.2.2), we have:
@@
all terms containing $x_{c}$ are commutators with a c-number, which therefore vanish. Thus, the Agarwal dissipator preserves its form under $\hat{T}(x_{c})$; hence, in contrast with the Lindblad dissipator, adjustment to the actual position $x_{c}$ of the harmonic oscillator is not required.